OpenStudy (anonymous):

Elimination 4. 4x + 3y = -45.5 -5x + 8y = -43

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

20x+15y=-227.5 -20x+32y=-172 ----------------------- 0x+47y=-399.5 y=-399.5/47=8.5 now substitute into one of original equations to get x